# An Adaptive Tangent Feature Perspective of Neural Networks

### 28 Dec 2023 v2

@@ Abstract
In order to better understand feature learning in neural networks, we propose and study linear models in tangent feature space where the features are allowed to be transformed during training. We consider linear feature transformations, resulting in a joint optimization over parameters and transformations with a bilinear interpolation constraint. We show that this relaxed optimization problem has an equivalent linearly constrained optimization with structured regularization that encourages approximately low rank solutions. Specializing to structures arising in neural networks, we gain insights into how the features and thus the kernel function change, providing additional nuance to the phenomenon of kernel alignment when the target function is poorly represented by tangent features. In addition to verifying our theoretical observations in real neural networks on a simple regression problem, we empirically show that an adaptive feature implementation of tangent feature classification has an order of magnitude lower sample complexity than the fixed tangent feature model on MNIST and CIFAR-10.
@@ 1 Introduction
  • We empirically evaluate our adaptive feature model in neural networks on MNIST and CIFAR-10, which provides an order of magnitude better sample complexity compared to fixed tangent features.
Our work is quite far from an exact characterization of real neural networks, and our empirical results indicate that adaptivity as we propose cannot fully explain the performance of real neural networks. Nevertheless, our framework extends the class of phenomena can be explained by linear models built on tangent features, and so we believe it to be a valuable contribution towards understanding neural networks.
@@
All proofs can be found in Appendix B.
@@ 3.1 Structureless feature learning
In order to go beyond ridgeless regression with feature learning, it is necessary to further constrain the structure of $\mathbf{M}$. By restricting $\mathbf{M}$ to operate independently on separate subspaces, the result is an effective group sparse penalty over the subspaces. As an extreme example that we detail in Section B.1, if $\mathbf{M}$ is constrained to be diagonal, operating independently on $P$ orthogonal subspaces of dimension 1, then the result is that $\widetilde{\Omega}$ is a conventional sparsity-inducing penalty. Concretely, we have $\widetilde{\Omega}=\|\cdot\|_{1}$ in the special case of $\Omega(\mathbf{M})=\|\mathbf{M}\|_{F}^{2}$ with a diagonal constraint.
@@ 3.2 Neural network structure
Figure 1: A more difficult task yield higher label kernel alignment. We perform regression using a multi-layer perceptron on 500 MNIST digits from classes 2 and 3. We construct target labels $y_{i}$ as the best linear fit of binary ±$\pm 1$ labels using random neural network features. Then we train two networks, one (left) trained to predict $y_{i}$, and one (right) trained to predict $\mathrm{sign}(y_{i})$. We present the adapted kernel and label kernel matrices for data points ordered according to $y_{i}$ and report the cosine similarity of the adapted kernel and the label kernel. The harder task of regression with binarized labels has a higher label kernel alignment. Further details are given in Section C.1.
@@
It is straightforward to verify that if $\omega_{1}$ and $\omega_{2}$ are strictly quasi-convex such that $\omega_{1}(1)=\omega_{2}(1)=0$, then $\omega_{1}\oplus\omega_{2}$ has the same properties; see Section B.2. We also take this opportunity to define the scalar effective penalty and its induced effective penalty, which will simplify notation:
@@ 4 Discussion
In this work, we have proposed a framework of adaptive feature learning using tangent features that sheds light on some phenomena observed in real neural networks. The most important question one should ask, however, is the degree to which real neural networks actually fit this model, and whether there still remains a gap in our understanding.
In terms of practical performance, we have found that the adaptive feature learning model offers significant sample complexity advantages over fixed neural tangent kernel features. In Figure 4, we compare the two models and real neural networks on MNIST and CIFAR-10. For the same test error, the adaptive feature model requires an order of magnitude fewer samples. However, given enough samples, both the adaptive and fixed feature models appear to converge to the same error floor, while the real neural networks significantly outperform both models and appear to have a faster convergence rate in the case of CNNs on CIFAR-10. This suggests that there are fundamental limitations of the tangent kernel feature space that even adaptive feature learning cannot avoid, although adaptivity does close a non-negligible part of the gap in the low-sample regime. An interesting direction for future work is to analyze the convergence rate of adaptive feature learning under Model A and determine under what conditions there might be improved rates over fixed tangent features.
Though a gap still exists between our model and real neural networks, it has nevertheless provided valuable insight into feature learning in neural networks. In addition to the previous dive into the effects on the adapted kernel, we provide additional discussion in Appendix A where we consider the difference between average tangent features and final tangent features, the connection between approximate low rank optimization and optimal benign interpolation in sparse settings, justification for the recently successful low rank adaptation (LoRA) method [37], and the effects of architecture on adaptive feature learning.
Figure 4: Adaptive feature learning improves low-sample performance. We perform 10-class classification on MNIST (left) and CIFAR-10 (right) using three different models: a linear model using fixed tangent features (blue, solid), the adaptive feature optimization in eq. 4 under Model A (orange, dashed), and standard neural networks (green, dash–dot). The adaptive feature model achieves the same performance as the non-adaptive tangent feature model with an order of magnitude fewer samples. Further details are given in Section C.2.
Acknowledgements
The experiments in this paper were written in PyTorch [38] and JAX [39] with code assistance from GitHub Copilot and ChatGPT. DL was supported by ARO grant 2003514594. SA and computing resources were supported by Richard G. Baraniuk via NSF grants CCF-1911094, IIS-1838177, and IIS-1730574; ONR grants N00014-18-1-2571, N00014-20-1-2534, and MURI N00014-20-1-2787; AFOSR grant FA9550-22-1-0060; and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.
Appendix A Additional Discussion
Here we collection additional discussion points not included in the main paper.
@@ Appendix A Additional Discussion
Appendix B Theoretical technical details
B.1 Diagonal $\mathbf{M}$ example
@@
B.2 Penalty function results
B.2.1 Joint penalty quasi-convexity
@@ B.2.2 Proof of Proposition 3
B.3 Proof of main results
@@ B.3.2 Proof of Theorem 2
Appendix C Experimental details
C.1 MNIST regression experiment
@@
C.2 MNIST and CIFAR-10 classification experiment
We report average test error with standard deviation error bars over 10 trials. Our models are a 5-layer multi-layer perceptron (MLP) with 128 units at each layer and ReLU activation before the final 10 outputs, implemented in JAX [39]. For layer biases, when applying $\widehat{\mathbf{M}}_{\ell}^{(2)}$, we also map $\mathbf{b}_{\ell}\mapsto\widehat{\mathbf{M}}_{\ell}^{(2)\mathsf{T}}\mathbf{b}_{\ell}$. For the feature models, we do not impose any explicit regularization other than random initialization. For the adaptive feature model, we initialize $\mathbf{M}_{\ell}^{(1)}=\mathbf{I}_{P_{\ell}}$ and $\mathbf{M}_{\ell}^{(2)}=\mathbf{I}_{Q_{\ell}}$. For each trial, we select $N\in\{125,250,500,1000,2000,4000,8000,16000,32000,60000\}$ training points at random and train using SGD with 0.01 learning rate and 0.9 momentum.
For CIFAR-10, we proceed similarly to MNIST, but instead we use a convolutional neural network (CNN) with 4 $3\times 3$ convolutional layers with 2×2$\times$2 × stride and average pooling after ReLU activation, doubling the number of channels each layer after first mapping from 3 to 16. We then flatten and pass through 2 dense layers with 128 units before the final 10 outputs. For convolutional kernels we use three $\mathbf{M}_{\ell}^{(k)}$ transformations, one for the flattened $3\times 3$ convolutional kernels, one for the input channels, and one for the output channels, which we also apply to the biases. We select $N\in\{125,250,500,1000,2000,4000,8000,16000,32000,50000\}$ training points at random with no data augmentation and train using the Adam optimizer [41] with 0.001 learning rate.

### 20 Feb 2024 v3

@@ Abstract
In order to better understand feature learning in neural networks, we propose and study linear models in tangent feature space where the features are allowed to be transformed during training. We consider linear feature transformations, resulting in a joint optimization over parameters and transformations with a bilinear interpolation constraint. We show that this relaxed optimization problem has an equivalent linearly constrained optimization with structured regularization that encourages approximately low rank solutions. Specializing to structures arising in neural networks, we gain insights into how the features and thus the kernel function change, providing additional nuance to the phenomenon of kernel alignment when the target function is poorly represented by tangent features. We verify our theoretical observations in the kernel alignment of real neural networks.
@@ 1 Introduction
Our work is quite far from an exact characterization of real neural networks. Nevertheless, our framework extends the class of phenomena can be explained by linear models built on tangent features, and so we believe it to be a valuable contribution towards understanding neural networks.
@@
All proofs can be found in Appendix A.
@@ 3.1 Structureless feature learning
In order to go beyond ridgeless regression with feature learning, it is necessary to further constrain the structure of $\mathbf{M}$. By restricting $\mathbf{M}$ to operate independently on separate subspaces, the result is an effective group sparse penalty over the subspaces. As an extreme example that we detail in Section A.1, if $\mathbf{M}$ is constrained to be diagonal, operating independently on $P$ orthogonal subspaces of dimension 1, then the result is that $\widetilde{\Omega}$ is a conventional sparsity-inducing penalty. Concretely, we have $\widetilde{\Omega}=\|\cdot\|_{1}$ in the special case of $\Omega(\mathbf{M})=\|\mathbf{M}\|_{F}^{2}$ with a diagonal constraint.
@@ 3.2 Neural network structure
Figure 1: A more difficult task yield higher label kernel alignment. We perform regression using a multi-layer perceptron on 500 MNIST digits from classes 2 and 3. We construct target labels $y_{i}$ as the best linear fit of binary ±$\pm 1$ labels using random neural network features. Then we train two networks, one (left) trained to predict $y_{i}$, and one (right) trained to predict $\mathrm{sign}(y_{i})$. We present the adapted kernel and label kernel matrices for data points ordered according to $y_{i}$ and report the cosine similarity of the adapted kernel and the label kernel. The harder task of regression with binarized labels has a higher label kernel alignment. Further details are given in Section B.1.
@@
It is straightforward to verify that if $\omega_{1}$ and $\omega_{2}$ are strictly quasi-convex such that $\omega_{1}(1)=\omega_{2}(1)=0$, then $\omega_{1}\oplus\omega_{2}$ has the same properties; see Section A.2. We also take this opportunity to define the scalar effective penalty and its induced effective penalty, which will simplify notation:
@@ 4 Discussion
Although our framework is far from capturing all of the complexities of neural networks, we have been able to shed some light on kernel alignment to close our gap in understanding. We now discuss a few possibilities for future work and connections to other closely related literature.
@@
Acknowledgements
The experiments in this paper were written in PyTorch [39] with code assistance from GitHub Copilot. DL was supported by ARO grant 2003514594. SA and computing resources were supported by Richard G. Baraniuk via NSF grants CCF-1911094, IIS-1838177, and IIS-1730574; ONR grants N00014-18-1-2571, N00014-20-1-2534, and MURI N00014-20-1-2787; AFOSR grant FA9550-22-1-0060; and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.
Appendix A Theoretical technical details
A.1 Diagonal $\mathbf{M}$ example
@@
A.2 Penalty function results
A.2.1 Joint penalty quasi-convexity
@@ A.2.2 Proof of Proposition 3
A.3 Proof of main results
@@ A.3.2 Proof of Theorem 2
Appendix B Experimental details
B.1 MNIST regression experiment